Which of the following is the weakest acid?

A

Phenol (Ka=1.3×1010)

B

Hydrocyanic acid (Ka=4.9×1010)

C

Acetic acid (Ka=1.8×1010)

D

Benzoic acid (Ka=6.5×1010)

 
 
 

Grade:12

1 Answers

Adarsh Pal
3 years ago
Which ofthe following is the weakest acid ? (1) Phenol(K = 1.3x10-10)(2) Hydrocyanic acid (K = 4.9x10-19)(3) Acetic acid (K = 1.8 x10-5)(4) Benzoic acid (K = 6.5 x10-) The pH of 0.1 M monobasic acid is 4.50. The acidity constant(K) ofthe monobasic acid is-(2)1.0 x 10-5 (3)1.0 x104 (1)1.0 x 10-7 (4)1.0 x10-4 8. Which ofthe following is the strongest base ? 13 CH NHCH.